Mike, a salvation army bell ringer, has 5 fewer quarters than nickels in his cup. If Mike has $5.95 in quarters and nickels, how many nickels does he have?
I need help extracting the algebraic equation from this story problem.

Answer by ptaylor(1894)

Let x=number of nickels in Mike's cup
Then x-5=number of quarters in Mike's cup
Let's now deal in pennies to make life a little more simple.
5x=amount of money that Mike has from the nickels
25(x-5)=amount of money that Mike has from the quarters
We are told that this adds up to 595. So our algebraic equation is:
5x+25(x-5)=595 get rid of parens (Now as you well know, if we want to deal in dollars, we can just as easily write this equation as:
0.05x+0.25(x-5)=5.95), but lets continue to deal in pennies
5x+25x-125=595 add 125 to both sides
5x+25x-125+125=595+125 collect like terms
30x=720 divide both sides by 30
x=24------------------------------------number of nickels
x-5=24-5=19---------------------------------number of quarters
CK
24(0.05)+19(0.25)=5.95
1.20+4.75=5.95
5.95=5.95

Hope this helps----ptaylor